OpenStudy (anonymous):

Find the equation of the line whose slope is 0 and that contains the point (-4, 2). Answer y + 4 = 0 x + 4 = 0 x - 2 = 0 y - 2 = 0

OpenStudy (hoblos):

slope is 0 so it is a horizontal line passing by (-4, 2) its equation is y=2-----> y - 2 = 0

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

my pleasure :D

OpenStudy (anonymous):

Always go back to y eeeequaaals mx + b y=mx+b (x,y): your point m: your slope b: y-intercept To approach this, you have (x,y), you have m, but you don't have b. You plug in the values you know and solve for b. 2 = (0)(-4) +b b=2 So we have our b-value, along with our m-value. We don't need the (x,y) to make our eqn of a line. Plug all of these in and we get our line. y = (0)*x +2 y = 2 Or like they put it, you can rearrange and get: y -2 = 0

OpenStudy (hoblos):

yup this works at any time ;)